# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@
Abstract.
@@
2020 Mathematics Subject Classification:
Primary 53B20, Secondary 32V05, 53C18
1. Introduction
The Fefferman construction in CR geometry associates a pseudo-Riemannian conformal manifold $(\mathcal{C},[g^{\mathrm{F}}])$, called the Fefferman space, to any non-degenerate CR manifold $(M,T^{1,0}M)$ ([11, 18]). The manifold $\mathcal{C}$ is a circle bundle over $M$ and the Fefferman metric $g^{\mathrm{F}}$ is an $S^{1}$-invariant metric on $\mathcal{C}$ which is conformally flat if and only if $M$ is CR isomorphic to the hyperquadric, i.e., the flat model of CR manifold. Since the conformal structure $[g^{\mathrm{F}}]$ is canonically defined from the CR structure, the construction enables us to apply conformal geometry methods to the study of CR manifolds.
The aim of this paper is to describe explicitly the Fefferman metric for a special class of CR $5$-manifolds, namely twistor CR manifolds, and apply it to $3$-dimensional conformal geometry. Given a conformal $3$-manifold $(\Sigma^{3},[g])$ of Riemannian signature, LeBrun’s twistor CR manifold ([17]) is defined as the space of projectivized complex null (co)vectors:
$M^{5}:=\{[\zeta]\in\mathbb{P}(\mathbb{C}T^{*}\Sigma)\ |\ g(\zeta,\zeta)=0\}.$
@@ 1. Introduction
$\widehat{\mathcal{C}}:=\{\zeta\in\mathbb{C}T^{*}\Sigma\ |\ g(\zeta,\zeta)=0,\,g(\zeta,\overline{\zeta})=1\}$
over $M$ can be identified with a double covering of the Fefferman space $\mathcal{C}$ by constructing an ‘adapted’ coframe $\{\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\}$ on $\widehat{\mathcal{C}}$ out of the Riemannian structure $g$. Here, $\theta$ is a contact form on $M$ determined by $g$, and $(\tilde{\theta}^{1},\tilde{\theta}^{2})$ is a ‘twisted’ coframe for $T^{1,0}M$ while $\tilde{\theta}^{3}$ restricts to a coframe for the $S^{1}$-fibers. In this coframe, (the pull-back of) the Fefferman metric $g^{\mathrm{F}}$ on $\widehat{\mathcal{C}}$ is written as
$g^{\mathrm{F}}=2(\tilde{\theta}^{1}\cdot\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\cdot\tilde{\theta}^{\overline{1}}+\theta\cdot\tilde{\theta}^{3}).$
Using this expression, we compute the Levi-Civita connection and the curvature tensor of $g^{\mathrm{F}}$ in terms of those of $g$. In particular, the Weyl curvature of $g^{\mathrm{F}}$ is described by the Hodge dual of the Cotton tensor of $g$. As a consequence, we can recover [19, Theorem 6.7], which states that the Fefferman metric is conformally flat if and only if $(\Sigma,[g])$ is conformally flat (Theorem 4.8).
As another application, we establish a correspondence between distinguished curves on $(M,T^{1,0}M)$ and $(\Sigma,[g])$.
Since null geodesics of $g^{\mathrm{F}}$ are conformally invariant up to reparametrizations, their projections define a CR invariant family of curves on $M$, called (null) chains. The infinitesimal generator $K$ of the $S^{1}$-action on $\widehat{\mathcal{C}}$ is a Killing vector field, so $c:=g^{\mathrm{F}}(K,\dot{\gamma}(t))$ is constant in $t$ for any null geodesic $\gamma$. The projection of $\gamma$ is called a chain if $c\neq 0$ and a null chain if $c=0$. Chains are transverse to the contact distribution $H=\mathrm{Re}\,T^{1,0}M$ and uniquely determined by the initial velocity. Thus, they can be considered as ‘geodesics’ in CR geometry. In contrast, null chains are tangent to $H$ and we also need information of the initial acceleration to specify a null chain; see [13].
@@
where $P$ denotes the Schouten tensor of $g\in[g]$. Due to its conformal invariance, conformal geodesics play an important role in general relativity to examine the structure of spacetime; see, e.g., [12, 9]. Although the conformal geodesic equation specifies distinguished parametrizations of $x(t)$, called projective parameters ([1]), we consider conformal geodesics as unparametrized curves in this paper.
By computing the null geodesic equation for the Fefferman metric $g^{\mathrm{F}}$, we prove that these invariant curves on $M$ and $\Sigma$ are related as follows:
@@
    For any conformal geodesic $x(t)\ (a\leq t\leq b)$ on $\Sigma$ and $[\zeta_{0}]\in M_{x(a)}$, there exists a unique chain (resp. a null chain) through $[\zeta_{0}]$ which projects to $x(t)$ when $g(\eta_{0},\dot{x}(a))\neq 0$ (resp. $g(\eta_{0},\dot{x}(a))=0$), where $\eta_{0}:=i\zeta_{0}\times\overline{\zeta_{0}}$.
This theorem asserts that any (null) chain projects to a conformal geodesic, and conversely we can lift any conformal geodesic to a (null) chain. The lift is determined by a choice of the lift of the endpoint $x(a)$, but in fact there is a canonical choice, namely the lift with $\eta_{0}$ being proportional to the initial velocity $\dot{x}(a)$. The lift is described as follows: We fix a metric $g\in[g]$ and a (local) orientation of $\Sigma$. For any regular curve $x(t)$ on $\Sigma$, we have a complex null (co)vector field $\zeta(t)$, unique up to the $\mathrm{U}(1)$-actions, such that $(\sqrt{2}\,\mathrm{Re}\,\zeta,\sqrt{2}\,\mathrm{Im}\,\zeta)$ gives an oriented orthonormal basis of $\dot{x}(t)^{\perp}\subset T_{x(t)}\Sigma$. The projective class $[\zeta(t)]$ is independent of the choice of $g\in[g]$ and we have a canonical lift $\tilde{x}(t):=(x(t),[\zeta(t)])$ on $M$, which is transverse to the contact distribution. Then, we can prove that $\tilde{x}(t)$ is a chain if and only if $x(t)$ is a conformal geodesic (Theorem 4.12).
There is an alternative description of the Fefferman space $(\widehat{\mathcal{C}},g^{\mathrm{F}})$ and the canonical lift $\tilde{x}(t)$ in terms of the unit tangent sphere bundle $S\Sigma\subset T\Sigma$ and the oriented orthonormal frame bundle $\mathcal{P}$ over $(\Sigma,g)$: Their relations are given by the bundle isomorphisms
@@
where $\eta:=i\zeta\times\overline{\zeta}=2\,\mathrm{Re}\,\zeta\times\mathrm{Im}\,\zeta$. By using the canonical $1$-form and the principal connection $1$-form
$\phi=\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix},\quad\omega=\begin{pmatrix}0&\omega^{3}&-\omega^{2}\\ -\omega^{3}&0&\omega^{1}\\ \omega^{2}&-\omega^{1}&0\end{pmatrix}$
@@
$G^{\mathrm{F}}=2(\phi^{1}\cdot\omega^{1}+\phi^{2}\cdot\omega^{2}+\phi^{3}\cdot\omega^{3}).$
@@
(1.1) $\tilde{x}(t)=\Bigl{(}x(t),\frac{\dot{x}(t)}{|\dot{x}(t)|}\Bigr{)}.$
@@
$F(\tilde{x},\xi):=\frac{G^{\mathrm{F}}(s_{*}\xi,s_{*}\xi)}{G^{\mathrm{F}}(K,s_{*}\xi)}\quad(\tilde{x}\in U\subset S\Sigma,\ \xi\in T_{\tilde{x}}S\Sigma\setminus\mathrm{Ker}\,\theta)$
for each choice of a local section $s\colon U\to\mathcal{P}$. The Kropina metric is not defined along the contact distribution since $K\lrcorner\,G^{\mathrm{F}}$ descends to the contact form $\theta$. Using the canonical lift (1.1), we define a functional $\mathscr{F}$ on the space of regular curves on $\Sigma$ by
[…]
where $s$ is a local section of $\mathcal{P}\to S\Sigma$ on a neighborhood of the curve $\tilde{x}(t)$. If $x(t)$ is a conformal geodesic, then $\tilde{x}(t)$ becomes a chain and hence $x(t)$ is a critical curve of $\mathscr{F}$ under the variations fixing the lifts $\tilde{x}(a),\tilde{x}(b)$ of the endpoints. Conversely, if $x(t)$ is a critical curve of $\mathscr{F}$, the Kropina length functional is stationary at $\tilde{x}(t)$ with respect to the variations given by the lift of variations of $x(t)$. Combined with the fact that the variations of the Kropina length in the vertical directions always vanish (Proposition 5.7), this implies that $\tilde{x}(t)$ becomes a chain. Consequently, conformal geodesics are characterized as the critical curves of $\mathscr{F}$ under the variations fixing $\tilde{x}(a),\tilde{x}(b)$ (Theorem 5.8).
The functional $\mathscr{F}$ is closely related to the so-called total torsion functional, which we denote by $\mathscr{T}$. For a regular curve $x(t)\ (a\leq t\leq b)$ on $(\Sigma,g)$, we fix an isometry $\dot{x}(a)^{\perp}\cong\dot{x}(b)^{\perp}$ by choosing oriented orthonormal basis $A=(e_{1},e_{2},\dot{x}(a)/|\dot{x}(a)|)\in\mathcal{P}_{x(a)}$ and $B=(e^{\prime}_{1},e^{\prime}_{2},\dot{x}(b)/|\dot{x}(b)|)\in\mathcal{P}_{x(b)}$ at the endpoints. Then, $\mathscr{T}[x(t)]\in\mathbb{R}/2\pi\mathbb{Z}$ is defined to be the angle from $\nu(a)$ to $\nu(b)$, where $\nu(t)$ is a unit normal vector field along $x(t)$ which is parallel with respect to $\nabla^{\perp}$. The total torsion depends on the choice of $A,B$, but it is conformally invariant once we fix the conformal classes of $A$ and $B$; see Proposition 5.11. If we use $A=s(\tilde{x}(a)),B=s(\tilde{x}(b))$, then we have
[…]
@@
[…]
In summary, we have the following
@@
A regular curve $x(t)\ (a\leq t\leq b)$ on $(\Sigma^{3},g)$ is a conformal geodesic if and only if it is a critical curve of the total torsion functional $\mathscr{T}$ under the variations fixing $(x(a),\dot{x}(a)/|\dot{x}(a)|)$ and $(x(b),\dot{x}(b)/|\dot{x}(b)|)$ with common $A,B$.
@@
[…]
@@
This paper is organized as follows: In §2, we review the notion of CR structure and define (null) chains on CR manifolds via the Fefferman metric. In §3, we recall the definition of conformal geodesic and introduce LeBrun’s twistor CR manifold over conformal $3$-manifolds. Then, in §4, we give an explicit description of the Fefferman metric for twistor CR manifolds and compute its connection, curvature, and null geodesic equation to prove Theorem 1.1 (1). Here, we also prove that the canonical lift of a conformal geodesic becomes a chain, which gives a partial proof to Theorem 1.1 (2); the full proof will be given in §5.1 after we reinterpret our constructions in the sphere/frame bundle picture. In §5, we give a variational characterization for conformal geodesics in dimension three via the Kropina metric for chains, and discuss its relation to the total torsion functional.
Notations: We use tensor indices $\alpha,\beta\ (=1,\dots,n)$ for $T^{1,0}M$, $i,j,k,l,m\ (=1,2,3)$ for $\Sigma$, and $a,b,c,d\ (=0,1,2,\overline{1},\overline{2},3)$ for $\widehat{\mathcal{C}}$. We adopt Einstein’s summation convention throughout this paper. For 1-forms $\varphi,\psi$, we define the symmetric product by
$(\varphi\cdot\psi)(V,W):=\frac{1}{2}\bigl{(}\varphi(V)\psi(W)+\varphi(W)\psi(V)\bigr{)}.$
Acknowledgment This work was partially supported by JSPS KAKENHI Grant Number 22K13922.
2. CR manifolds and the Fefferman metric
2.1. CR structures
An almost CR structure on a $(2n+1)$-dimensional $C^{\infty}$ manifold $M$ is defined to be a rank $n$ complex vector sub-bundle $T^{1,0}M\subset\mathbb{C}TM$ satisfying $T^{1,0}M\cap T^{0,1}M=0$, where $T^{0,1}M:=\overline{T^{1,0}M}$. If the space of sections of $T^{1,0}M$ is closed under the Lie bracket, it is called a CR structure. Since $T^{1,0}M$ contains no non-zero real vector, the real part $H:=\mathrm{Re}\,T^{1,0}M\subset TM$ becomes a rank $2n$ real sub-bundle satisfying $\mathbb{C}H=T^{1,0}M\oplus T^{0,1}M$.

### v4

@@
[…]
The Fefferman Metric for Twistor CR Manifolds
[…]
The Fefferman Metric for Twistor CR Manifolds
and Conformal Geodesics in Dimension Three
[…]
Taiji MARUGAME
[…]
T. Marugame
[…]
Department of Mathematics, The University of Electro-Communications,
[…]
Received June 03, 2025, in final form October 16, 2025; Published online October 24, 2025
[…]
@@
[…]
53B20; 32V05; 53C18
1 Introduction
The Fefferman construction in CR geometry associates a pseudo-Riemannian conformal manifold $\bigl(\mathcal{C},\bigl[g^{\mathrm{F}}\bigr]\bigr)$, called the Fefferman space, to any non-degenerate CR manifold $\bigl(M,T^{1,0}M\bigr)$} [14, 22]. The manifold $\mathcal{C}$ is a circle bundle over $M$ and the Fefferman metric $g^{\mathrm{F}}$ is an $S^{1}$-invariant metric on $\mathcal{C}$ which is conformally flat if and only if $M$ is CR isomorphic to the hyperquadric, i.e., the flat model of CR manifold. Since the conformal structure $\bigl[g^{\mathrm{F}}\bigr]$ is canonically defined from the CR structure, the construction enables us to apply conformal geometry methods to the study of CR manifolds.
The aim of this paper is to describe explicitly the Fefferman metric for a special class of CR $5$-manifolds, namely twistor CR manifolds, and apply it to $3$-dimensional conformal geometry. Given a conformal $3$-manifold $\bigl(\Sigma^{3},[g]\bigr)$ of Riemannian signature, LeBrun’s twistor CR manifold [21] is defined as the space of projectivized complex null (co)vectors:
$M^{5}:=\{[\zeta]\in\mathbb{P}(\mathbb{C}T^{*}\Sigma)\mid g(\zeta,\zeta)=0\}.$
@@ 1 Introduction
$\widehat{\mathcal{C}}:=\bigl\{\zeta\in\mathbb{C}T^{*}\Sigma\mid g(\zeta,\zeta)=0,\,g\bigl(\zeta,\overline{\zeta}\bigr)=1\bigr\}$
over $M$ can be identified with a double covering of the Fefferman space $\mathcal{C}$ by constructing an ‘adapted’ coframe $\bigl(\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\bigr)$ on $\widehat{\mathcal{C}}$ out of the Riemannian structure $g$. Here, $\theta$ is a contact form on $M$ determined by $g$, and $\bigl(\tilde{\theta}^{1},\tilde{\theta}^{2}\bigr)$ is a ‘twisted’ coframe for $T^{1,0}M$ while $\tilde{\theta}^{3}$ restricts to a coframe for the $S^{1}$-fibers. In this coframe, (the pullback of) the Fefferman metric $g^{\mathrm{F}}$ on $\widehat{\mathcal{C}}$ is written as
$g^{\mathrm{F}}=2\bigl(\tilde{\theta}^{1}\cdot\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\cdot\tilde{\theta}^{\overline{1}}+\theta\cdot\tilde{\theta}^{3}\bigr).$
Using this expression, we compute the Levi-Civita connection and the curvature tensor of $g^{\mathrm{F}}$ in terms of those of $g$. In particular, the Weyl curvature of $g^{\mathrm{F}}$ is described by the Hodge dual of the Cotton tensor of $g$. As a consequence, we can recover [23, Theorem 6.7], which states that the Fefferman metric is conformally flat if and only if $(\Sigma,[g])$ is conformally flat (Theorem 4.8).
As another application, we establish a correspondence between distinguished curves on the CR manifold $\bigl(M,T^{1,0}M\bigr)$ and the conformal manifold $(\Sigma,[g])$.
Since null geodesics of $g^{\mathrm{F}}$ are conformally invariant up to reparametrizations, their projections define a CR invariant family of curves on $M$, called ((null)) chains. The infinitesimal generator $K$ of the $S^{1}$-action on $\widehat{\mathcal{C}}$ is a Killing vector field, so $c:=g^{\mathrm{F}}(K,\dot{\gamma}(t))$ is constant in $t$ for any null geodesic $\gamma$. The projection of $\gamma$ is called a chain if $c\neq 0$ and a null chain if $c=0$. Chains are transverse to the contact distribution $H=\operatorname{Re}T^{1,0}M$ and uniquely determined by the initial velocity. Thus, they can be considered as ‘geodesics’ in CR geometry. In contrast, null chains are tangent to $H$ and we also need information of the initial acceleration to specify a null chain; see [17].
@@
where $P$ denotes the Schouten tensor of $g\in[g]$. Due to its conformal invariance, conformal geodesics play an important role in general relativity to examine the structure of spacetime; see, e.g., [12, 15]. Although the conformal geodesic equation specifies distinguished parametrizations of $x(t)$, called projective parameters [1], we consider conformal geodesics as unparametrized curves in this paper.
By computing the null geodesic equation for the Fefferman metric $g^{\mathrm{F}}$, we prove that these invariant curves on $M$ and $\Sigma$ are related as follows.
@@
    For any conformal geodesic $x(t)$, $a\leq t\leq b$, on $\Sigma$ and $\zeta_{0}]\in M_{x(a)}$, there exists a unique chain ((resp. a null chain)) through $\zeta_{0}]$ which projects to $x(t)$ when $g(\eta_{0},\dot{x}(a))\neq 0$ ((resp. $g(\eta_{0},\dot{x}(a))=0)$, where $\eta_{0}:={\rm i}\zeta_{0}\times\overline{\zeta_{0}}$.
This theorem asserts that any (null) chain projects to a conformal geodesic, and conversely we can lift any conformal geodesic to a (null) chain. The lift is determined by a choice of the lift of the endpoint $x(a)$, but in fact there is a canonical choice, namely the lift with $\eta_{0}$ being proportional to the initial velocity $\dot{x}(a)$. The lift is described as follows: We fix a metric $g\in[g]$ and a (local) orientation of $\Sigma$. For any regular curve $x(t)$ on $\Sigma$, we have a complex null (co)vector field $\zeta(t)$, unique up to the $\mathrm{U}(1)$-actions, such that $\bigl(\sqrt{2}\operatorname{Re}\zeta,\sqrt{2}\operatorname{Im}\zeta\bigr)$ gives an oriented orthonormal basis of $\dot{x}(t)^{\perp}\subset T_{x(t)}\Sigma$. The projective class $\zeta(t)]$ is independent of the choice of $g\in[g]$ and we have a canonical lift $\tilde{x}(t):=(x(t),[\zeta(t)])$ on $M$, which is transverse to the contact distribution. Then, we can prove that $\tilde{x}(t)$ is a chain if and only if $x(t)$ is a conformal geodesic (Theorem 4.12).
There is an alternative description of the Fefferman space $\bigl(\widehat{\mathcal{C}},g^{\mathrm{F}}\bigr)$ and the canonical lift $\tilde{x}(t)$ in terms of the unit tangent sphere bundle $S\Sigma\subset T\Sigma$ and the oriented orthonormal frame bundle $\mathcal{P}$ over $(\Sigma,g)$: Their relations are given by the bundle isomorphisms
@@
where we define $\eta:={\rm i}\zeta\times\overline{\zeta}=2\operatorname{Re}\zeta\times\operatorname{Im}\zeta$ with the cross product with respect to the representative metric $g$. By using the canonical $1$-form and the principal connection $1$-form
$\phi=\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix},\qquad\omega=\begin{pmatrix}\hphantom{-}0&\hphantom{-}\omega^{3}&-\omega^{2}\\ -\omega^{3}&\hphantom{-}0&\hphantom{-}\omega^{1}\\ \hphantom{-}\omega^{2}&-\omega^{1}&\hphantom{-}0\end{pmatrix}$
@@
$G^{\mathrm{F}}=2\bigl(\phi^{1}\cdot\omega^{1}+\phi^{2}\cdot\omega^{2}+\phi^{3}\cdot\omega^{3}\bigr).$
@@
$\tilde{x}(t)=\biggl(x(t),\frac{\dot{x}(t)}{|\dot{x}(t)|}\biggr).$ (1.1)
@@
$F(\tilde{x},\xi):=\frac{G^{\mathrm{F}}(s_{*}\xi,s_{*}\xi)}{G^{\mathrm{F}}(K,s_{*}\xi)},\qquad\tilde{x}\in U\subset S\Sigma,\quad\xi\in T_{\tilde{x}}S\Sigma\setminus\operatorname{Ker}\theta,$
for each choice of a local section $s\colon U\to\mathcal{P}$. The Kropina metric is not defined along the contact distribution since $K\lrcorner\,G^{\mathrm{F}}$ descends to the contact form $\theta$. Using the canonical lift (1.1), we define a functional $\mathscr{L}$ on the space of regular curves on $\Sigma$ by
$\mathscr{L}[x(t)]:=\int_{a}^{b}F\bigl(\tilde{x}(t),\dot{\tilde{x}}(t)\bigr)\,{\rm d}t\biggl(=2\int_{a}^{b}\omega^{3}\bigl(s_{*}\dot{\tilde{x}}\bigr)\,{\rm d}t\biggr),$
where $s$ is a local section of $\mathcal{P}\to S\Sigma$ on a neighborhood of the curve $\tilde{x}(t)$. If $x(t)$ is a conformal geodesic, then $\tilde{x}(t)$ becomes a chain and hence $x(t)$ is a critical curve of $\mathscr{L}$ under the variations fixing the lifts $\tilde{x}(a)$, $\tilde{x}(b)$ of the endpoints. Conversely, if $x(t)$ is a critical curve of $\mathscr{L}$, the Kropina length functional is stationary at $\tilde{x}(t)$ with respect to the variations given by the lift of variations of $x(t)$. Combined with the fact that the variations of the Kropina length in the vertical directions always vanish (Proposition 5.7), this implies that $\tilde{x}(t)$ becomes a chain. Consequently, conformal geodesics are characterized as the critical curves of $\mathscr{L}$ under the variations fixing $\tilde{x}(a)$, $\tilde{x}(b)$ (Theorem 5.8).
The functional $\mathscr{L}$ is closely related to the so-called total torsion functional, which we denote by $\mathscr{T}$. For a regular curve $x(t)$, $a\leq t\leq b$, on $(\Sigma,g)$, we fix an isometry $\dot{x}(a)^{\perp}\cong\dot{x}(b)^{\perp}$ by choosing oriented orthonormal basis $A=(e_{1},e_{2},\dot{x}(a)/|\dot{x}(a)|)\in\mathcal{P}_{x(a)}$ and $B=\bigl(e^{\prime}_{1},e^{\prime}_{2},\dot{x}(b)/|\dot{x}(b)|\bigr)\in\mathcal{P}_{x(b)}$ at the endpoints. Then, $\mathscr{T}[x(t)]\in\mathbb{R}/2\pi\mathbb{Z}$ is defined to be the angle from $\nu(a)$ to $\nu(b)$, where $\nu(t)$ is a unit normal vector field along $x(t)$ which is parallel with respect to $\nabla^{\perp}$. The total torsion depends on the choice of $A$, $B$, but it is conformally invariant once we fix the conformal classes of $A$ and $B$; see Proposition 5.11. If we use $A=s(\tilde{x}(a)),B=s(\tilde{x}(b))$, then we have
$\mathscr{L}[x(t)]\equiv 2\mathscr{T}[x(t)]\mod 2\pi\mathbb{Z}$
@@
$\mathscr{T}[x(t)]\equiv-\int_{a}^{b}\tau(t)|\dot{x}(t)|\,{\rm d}t+c\mod 2\pi\mathbb{Z}.$
In summary, we have the following.
@@
A regular curve $x(t)$, $a\leq t\leq b$, on $\bigl(\Sigma^{3},g\bigr)$ is a conformal geodesic if and only if it is a critical curve of the total torsion functional $\mathscr{T}$ under the variations fixing $(x(a),\dot{x}(a)/|\dot{x}(a)|)$ and $(x(b),\dot{x}(b)/|\dot{x}(b)|)$ with common $A$, $B$.
@@
$\int_{a}^{b}\tau(t)|\dot{x}(t)|\,{\rm d}t=\int_{a}^{b}\frac{\det(\dot{x},\nabla_{\dot{x}}\dot{x},\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x})}{|\dot{x}\times\nabla_{\dot{x}}\dot{x}|^{2}}|\dot{x}|\,{\rm d}t.$
@@
Finally, we comment on the relation of our results to the general theory of Cartan geometry. Conformal geometry and CR geometry are both examples of so-called parabolic geometry, that is, Cartan geometry modeled on a homogeneous space $G/P$ where $G$ is a semi-simple Lie group and $P$ is a parabolic subgroup. Following the pioneering works of Cartan and Tanaka, the general theory of parabolic geometry has been extensively developed; see [7] for an exposition of the theory together with historical remarks. In particular, Fefferman’s construction in CR geometry has been generalized within the broader framework of extensions $G\subset\widetilde{G}$ of symmetry groups [6, 7]. The twistor CR manifold can also be reinterpreted in this setting. In fact, LeBrun’s twistor CR structure corresponds to the special case ($p=q=0$) of the twistor construction of CR structure from quaternionic contact structures associated with the extension $\mathrm{Sp}(p+1,q+1)\subset\mathrm{SU}(2p+2,2q+2)$ described in [7, Section 4.5.5]. We also remark that Sato–Yamaguchi [25] introduced the twistor CR structure on the tangent sphere bundle over conformal 3-manifolds by a Cartan geometric method independently of LeBrun. Moreover, a general theory of distinguished families of curves in parabolic geometry, including CR chains and conformal geodesics, has been developed by Čap–Slovák–Žádník [8]. However, the correspondence between distinguished curves in the general setting of Fefferman-type constructions does not yet appear to be fully understood. Developing a general theory in this direction would therefore be of considerable interest, and the present paper may be viewed as a case study illustrating the correspondence between distinguished curves and their variational character.
This paper is organized as follows: In Section 2, we review the notion of CR structure and define (null) chains on CR manifolds via the Fefferman metric. In Section 3, we recall the definition of conformal geodesic and introduce LeBrun’s twistor CR manifold over conformal $3$-manifolds. Then, in Section 4, we give an explicit description of the Fefferman metric for twistor CR manifolds and compute its connection, curvature, and null geodesic equation to prove Theorem 1.1 (1). Here, we also prove that the canonical lift of a conformal geodesic becomes a chain, which gives a partial proof to Theorem 1.1 (2); the full proof will be given in Section 5.1 after we reinterpret our constructions in the sphere/frame bundle picture. In Section 5, we give a variational characterization for conformal geodesics in dimension three via the Kropina metric for chains, and discuss its relation to the total torsion functional.
Notation.
We use tensor indices $\alpha,\beta=1,\dots,n$ for $T^{1,0}M$, $i,j,k,l,m=1,2,3$ for $\Sigma$, and $a,b,c,d=0,1,2,\overline{1},\overline{2},3$ for $\widehat{\mathcal{C}}$. We adopt Einstein’s summation convention throughout this paper. For 1-forms $\varphi$, $\psi$, we define the symmetric product by
$(\varphi\cdot\psi)(V,W):=\frac{1}{2}(\varphi(V)\psi(W)+\varphi(W)\psi(V)).$
2 CR manifolds and the Fefferman metric
2.1 CR structures
An almost CR structure on a $(2n+1)$-dimensional $C^{\infty}$ manifold $M$ is defined to be a rank $n$ complex vector sub-bundle $T^{1,0}M\subset\mathbb{C}TM$ satisfying $T^{1,0}M\cap T^{0,1}M=0$, where $T^{0,1}M:=\overline{T^{1,0}M}$}. If the space of sections of $T^{1,0}M$ is closed under the Lie bracket, it is called a CR structure. Since $T^{1,0}M$ contains no non-zero real vector, the real part $H:=\operatorname{Re}T^{1,0}M\subset TM$ becomes a rank $2n$ real sub-bundle satisfying $\mathbb{C}H=T^{1,0}M\oplus T^{0,1}M$.